# DreamSim: Learning New Dimensions of Human Visual Similarity using Synthetic Data

### v1

@@ 1 Introduction
In this paper, we introduce a new perceptual metric, which bridges the gap between lower-level patch-based metrics and broad categorical comparisons. We collect a new dataset named NIGHTS – Novel Image Generations with Human-Tested Similarity – containing human similarity judgments over image triplets. Each triplet consists of a reference image and two perturbed versions, along with human judgments as to which version is most similar to the reference (Fig. 1). We use iterative filtering together with recent diffusion models to collect our dataset, which is designed to capture image sets that are cognitively impenetrable (i.e. result in consistent decisions across different individuals) yet showcase rich variations. For example, our data contains images of similar object appearance, viewing angles, camera poses, overall layout, etc. This data differs qualitatively from prior low-level datasets zhang2018unreasonable , which focused on perturbations like blurring and adding noise, and from previous high-level datasets hebart2019things ; hebart2023things , which showed variation just at the level of categories (e.g. “is an image of a kitchen more like an image of a giraffe or an image of a beach”).
@@ 2 Related work
Perceptual similarity. Classical metrics such as Manhattan $\ell_{1}$, Euclidean $\ell_{2}$, MSE, and PSNR use point-wise difference to measure similarity, thus failing to capture important joint image statistics. Patch-based metrics, including SSIM wang2004image , FSIM zhang2011fsim , and HDR-VDP-2 mantiuk2011hdr tackle this issue and are widely used in applications involving photo-metric distortions such as image quality assessment. However, they do not capture the nuances of human vision when more structural ambiguity is present sampat2009complex and are not suited for more complex image generation tasks.
With the deep learning revolution, classical metrics have been replaced by learning-based metrics gatys2015neural ; johnson2016perceptual ; dosovitskiy2016generating . These metrics are defined in the space of deep features extracted from pre-trained networks, such as VGG simonyan2014very or AlexNet alexnet . Amir and Weiss amir2021understanding demonstrate that even untrained networks can be adapted as perceptual metrics. Zhang et alzhang2018unreasonable observe that feature-based metrics outperform classical metrics, across different convolutional architectures and learning paradigms, suggesting that perceptual similarity was an emergent property in deep representations. Further tuning on the perceptual data yields improvements, such as in LPIPS zhang2018unreasonable , PIE-APP prashnani2018pieapp , or DPAM in the audio domain manocha2020differentiable ; manocha2021cdpam . Further improvements include ensembling for robustness kettunen2019lpips , antialiasing zhang2019making ; henaff2015geodesics ; ding2020image for stability, and global descriptors for texture ding2020image . Hebart et al.muttenthaler2022human provide insight into high-level human similarity by training on a subset of the THINGS hebart2019things dataset, focusing on concept similarity and omitting visual cues for images within a category.
While strong computer vision features make for strong perceptual metrics, counter-intuitively, they eventually become decorrelated with perceptual similarity dehghani2023scaling ; kumar2022better . Today, the predominantly-used perceptual metric is LPIPS, operating on 64$\times$64 patches.
@@ 3 Perceptual dataset collection
While previous datasets focus on low-level, patch-based distortionszhang2018unreasonable ; PONOMARENKO201557 or high-level, categorical hebart2019things distortions, we aim to close the gap, capturing distortions including mid-level changes. We aim to produce images with an underlying semantic commonality, but variations in a diversity of factors, such as style, color, pose, and other details, so that a human can assess their visual relationship. In Sec. 3.1, we describe our data generation pipeline – prompting Stable Diffusion for related images of a given category, leveraging its natural image prior for variations within the category. We then describe our mechanism for collecting cognitively impenetrable perceptual judgments in Sec. 3.2.
@@ 3.1 Generating images with varied distortions
We leverage Stable Diffusion rombach2022high , which generates diverse and high-quality images that adhere to a given text prompt. We sample images with a prompt of the same category, using the structure "An image of a <category>". The <category> is drawn from image labels in popular datasets – ImageNet deng2009imagenet , CIFAR-10 krizhevsky2009learning , CIFAR-100 krizhevsky2009learning , Oxford 102 Flower nilsback2008automated , Food-101 bossard2014food , and SUN397 xiao2010sun . While the ImageNet labels make up $\sim\!55\%$ of the categories, the superset of categories across these datasets encompass a wide range of objects and scenes.
As illustrated in Fig.1, the resulting images display mid-level variations, such as pose, perspective, and shape, falling in between pixel/patch-based distortions and categorical differences. Additional triplets and the full list of categories are in the Supplementary Materials (SM). We generate an initial set of 100,000 image triplets, which are later filtered during the labeling process. Each triplet consists of a reference image and two “distortions”, generated from the same prompt.
@@ 3.2 Human perceptual judgments
We strive to gather perceptual judgments that are automatic (requiring little to no cognition), stable (invariant to changes in mental representation), and shared across humans, which requires a cognitively impenetrable judging process. Our main dataset consists of two alternative forced choice (2AFC) judgments on a triplet. Additionally, we collect just noticeable difference (JND) judgments on image pairs, corroborating our findings. Table. 1 shows an overview of our dataset, compared to prior perceptual datasets.
@@
To collect cognitively impenetrable triplets, we design our 2AFC experiments to ensure that each triplet included in the dataset obtains a unanimous vote for either $\tilde{x}_{0}$ or $\tilde{x}_{1}$. We divide our experiments into 10 rounds, starting with 100,000 triplets, advancing instances that maintain unanimous votes. Triplets retained through all 10 rounds can earn a maximum of 10 votes, but may have less as a result of the aforementioned sentinels.
Ultimately, our 2AFC dataset $\mathcal{D}^{\text{2afc}}=\{(x,\tilde{x}_{0},\tilde{x}_{1}),y\}$ consists of 20,019 triplets with an average of 7 unanimous votes each. We partition our resulting dataset into train, validation, and test components with a random 80/10/10 split. We intend on releasing our dataset to the public.
@@
Intuitively, an image pair that falls below the JND threshold should be more likely to be selected as similar in the 2AFC test. To assess how well our perceptual tests agree, we collect judgments on the triplets from the 2AFC test set, filtering for triplets that “straddle” the JND threshold. Specifically, given triplet $(x,\tilde{x}_{0},\tilde{x}_{1})\in\mathcal{D}^{\text{2afc}}$, we independently collect JND judgments on pairs $(x,\tilde{x}_{0})$ and $(x,\tilde{x}_{1})$, keeping triplets where humans find one and only one to be identical. In total, our JND dataset contains 608 triplets $\mathcal{D}^{\text{jnd}}=\{(x,\tilde{x}_{0},\tilde{x}_{1}),s\}$, where users chose one of $s\in\{0,1\}$ to be identical.
@@ 4.1 Embeddings as a distance metric
We denote a distance between two images as $D(\cdot,\cdot\hskip 1.42262pt;f_{\theta})$, where $f_{\theta}$ is a feature extractor. We evaluate a variety of state-of-the-art candidate embeddings and metrics. LPIPS zhang2018unreasonable and DISTS ding2020image use CNN backbones, whereas DINO caron2021emerging , CLIP radford2021learning , OpenCLIP Cherti2022ReproducibleSL , and MAE he2021masked use transformer-based backbones dosovitskiy2020image . Following standard practice, distance $D(x,\tilde{x};f_{\theta})=1-\text{cos}\big{(}f_{\theta}(x),f_{\theta}(\tilde{x})\big{)}$ is taken as the cosine distance between the CLS tokens taken from the last layer for DINO and MAE (before and after the layer normalization, respectively), and the embedding vector for CLIP and OpenCLIP. For LPIPS and DISTS, $D(x,\tilde{x};f_{\theta})$ is simply the distance metric itself. Given a triplet $(x,\tilde{x}_{0},\tilde{x}_{1})$, and a feature extractor $f_{\theta}$, the vote is calculated as:
@@ 4.2 Learning an improved metric
Tuning. Next, we investigate the best method to tune or modify feature extractor $f_{\theta}$. A challenge is the large number of parameters (billions, for ViT backbones), compared to the relatively small amount of perceptual data. Inspired by zhang2018unreasonable , we first try tuning via an MLP head, adding a 1-hidden-layer MLP with a residual connection on top of the pre-trained embeddings. We compare this to fine-tuning through the pre-trained backbones using Low-Rank Adaptation (LoRA) hu2021lora , which, in preliminary experiments, achieved better results than full fine-tuning. Using LoRA ($r=16$, dropout $p=0.3$, $\alpha=0.5$) we tune approximately 0.67% of each model’s parameters. In all experiments, LoRA significantly outperformed using the MLP head.
@@ 5.1 How well do existing metrics align with human judgments?
In Figure 5, we show how well metrics align with humans on our dataset, across low-level metrics (PSNR, SSIM), prior learned metrics (LPIPS, DIST) and recent large vision model embeddings (MAE, CLIP, DINO). The best-performing configurations of existing base models are DINO B/16, MAE B/16, CLIP B/32, and OpenCLIP B/32 (additional settings are in the SM). However, existing metrics have significant misalignments with humans; for example, DINO has approximately 10% disagreement. To improve performance, we finetune with LoRA (as described in Section 4.2), denoted as Tuned - LoRA, achieving significant alignment improvements on pre-trained baselines (Fig. 5) over training an MLP head (Tuned - MLP). Ensembling models by concatenating OpenCLIP, CLIP, and DINO features achieves the highest score across baselines (90.8%), MLP-tuned models (93.4%), and LoRA-tuned models (96.2%). We refer to the best LoRA-tuned ensemble model as DreamSim, and use it for subsequent analysis and applications.
Do different ways of measuring human perception agree? We corroborate our results on our 2AFC test set by evaluating the models against JND perceptual judgments. The high correlation between 2AFC and JND scores indicates that our 2AFC judgments capture a general sense of similarity that can replicated across different settings (Fig. 5).
Does improved mid-level perception lead to improved low-level and high-level perception? We use the same models as in Fig. 5 to examine how performance on our dataset corresponds to performance on two other datasets: BAPPS, containing image triplets with low-level augmentations, and THINGS, containing triplets with categorical variations. Training on our dataset improves human alignment on BAPPS but decreases alignment on THINGS, suggesting that humans, when making an automatic judgment, tend to focus on appearance-based visual differences (captured in BAPPS and our dataset) rather than high-level object category. For detailed results, refer to the SM.
@@ 5.2 What image attributes affect similarity decisions?
Figure 7: Alignment with semantic and low-level metrics. We select random triplets from the COCO dataset and assess how well low-level and semantic metrics predict model judgments. Compared to LPIPS & DISTS, our model is more sensitive to objects that appear in individual instances ("things") and the presence of people, and less sensitive to "stuff" categories, such as sky and road. Compared to OpenCLIP & DINO, color explains more of our model’s decisions. We mark our model’s results with a dashed pink line for comparison.
@@
Sensitivity to image ablations. We study which image properties affect our model’s decisions by ablating properties of the image triplets, such as orientation, color, shading, structure, and foreground or background components. To do this, we modify the triplets to change or discard each attribute and compute the agreement between the model decisions on the modified and unmodified triplets. This operation measures the model’s robustness when a given attribute is missing. As shown in Fig. 6, our model is least affected by orientation (tested by flipping the reference image), but ablating color or shading (the L or AB channels in LAB color space), structure  gatys2015texture , or color and shading together with contours chan2022learning has larger impacts. We conclude that color, shading, and structure contain critical information for our model, but our model can tolerate differences in orientation. To ablate foreground and background components, we segment the image carvekit and replace the corresponding pixels with random uniform noise. This operation removes the color, shading, and texture of the region, even though the same outline remains in both. Removing the background has a smaller impact on model alignment compared to removing the foreground, suggesting that the foreground color and texture is more important for similarity recognition in our model compared to the background.
@@
Alignment with semantic and low-level metrics. Given that some image attributes, like semantics, are difficult to ablate, we complement our ablation analysis by examining how well semantic and handcrafted low-level features align with model judgments. We use 10,000 random image triplets from MS-COCO lin2014microsoft annotated with individual object instances ("things") and background categories such as grass and sky ("stuff"). We compare our metric to LPIPS, DISTS, DINO, and OpenCLIP (Fig. 7). Our model aligns less strongly with RGB-color histogram similarity (calculated using histogram intersection swain1991color ) compared to LPIPS and DISTS, however is more sensitive to color compared to other large vision models, while none align with depth map similarity Eigen2014DepthMP ; Ranftl2020 . Next, we compute alignment with the things-category and stuff-category histograms, which summarize the area occupied by each semantic category. Our model aligns better with things-histogram similarity, whereas other models are comparatively more sensitive to stuff. This result suggests that our model is relatively more sensitive to foreground object instances. Finally, we compute how often metrics align with the per-category intersection of categorical histograms. Our model aligns best with the presence of people; additional results in the SM.
@@
Figure 8: Nearest-neighbor retrieval across different metrics. Nearest neighbor retrieval on ImageNet-R (top) and COCO (bottom), comparing different metrics. Although the datasets include images outside of our training domain, our model consistently retrieves neighbors with similar appearance and class to that of the query image. On ImageNet-R, our metric is preferred by users 36.8% of the time, followed by OpenCLIP (28.7%). Similarly, on COCO, users prefer our metric 33.6% of the time, with DINO being the second choice (28.1%).

### v3

@@ 1 Introduction
In this paper, we introduce a new perceptual metric, which bridges the gap between lower-level patch-based metrics and broad categorical comparisons. We collect a new dataset named NIGHTS – Novel Image Generations with Human-Tested Similarity – containing human similarity judgments over image triplets. Each triplet consists of a reference image and two perturbed versions, along with human judgments as to which version is most similar to the reference (Fig. 1). We use iterative filtering together with recent diffusion models to collect our dataset, which is designed to capture image sets that are cognitively impenetrable (i.e. result in consistent decisions across different individuals) yet showcase rich variations. For example, our data contains images of similar object appearance, viewing angles, camera poses, overall layout, etc. This dataset differs qualitatively from prior low-level datasets zhang2018unreasonable , which focused on perturbations like blurring and adding noise, and from previous high-level datasets hebart2019things ; hebart2023things , which showed variation just at the level of categories (e.g. “is an image of a kitchen more like an image of a giraffe or an image of a beach”).
@@ 2 Related work
Perceptual similarity. Classical metrics such as Manhattan $\ell_{1}$, Euclidean $\ell_{2}$, MSE, and PSNR use point-wise difference to measure similarity, thus failing to capture important joint image statistics. Patch-based metrics, including SSIM wang2004image , FSIM zhang2011fsim , and HDR-VDP-2 mantiuk2011hdr tackle this issue and are widely used in applications involving photometric distortions such as image quality assessment. However, they do not capture the nuances of human vision when more structural ambiguity is present sampat2009complex and are not suited for more complex image generation tasks.
With the deep learning revolution, classical metrics have been replaced by learning-based metrics gatys2015neural ; johnson2016perceptual ; dosovitskiy2016generating . These metrics are defined in the space of deep features extracted from pre-trained networks, such as VGG simonyan2014very or AlexNet alexnet . Amir and Weiss amir2021understanding demonstrate that even untrained networks can be adapted as perceptual metrics. Zhang et alzhang2018unreasonable observe that feature-based metrics outperform classical metrics across different convolutional architectures and learning paradigms, suggesting that perceptual similarity is an emergent property in deep representations. Further tuning on the perceptual data yields improvements, such as in LPIPS zhang2018unreasonable , PIE-APP prashnani2018pieapp , or DPAM in the audio domain manocha2020differentiable ; manocha2021cdpam . Further improvements include ensembling for robustness kettunen2019lpips , antialiasing for stability zhang2019making ; henaff2015geodesics ; ding2020image , and global descriptors for texture ding2020image . Muttenthaler et al.muttenthaler2022human provide insight into high-level human similarity by training on a subset of the THINGS hebart2019things dataset, focusing on concept similarity and omitting visual cues for images within a category.
While strong computer vision features make for strong perceptual metrics, counterintuitively, they eventually become decorrelated with perceptual similarity dehghani2023scaling ; kumar2022better . Today, the predominantly-used perceptual metric is LPIPS, operating on 64$\times$64 patches.
@@ 3 Perceptual dataset collection
While previous datasets focus on low-level, patch-based distortions zhang2018unreasonable ; PONOMARENKO201557 or high-level, categorical hebart2019things changes, we aim to close the gap, capturing distortions including mid-level variations. We aim to produce images with an underlying semantic commonality, but variations in a diversity of factors, such as style, color, pose, and other details, so that a human can assess their visual relationship. In Section 3.1, we describe our data generation pipeline – we prompt Stable Diffusion for related images of a given category, leveraging its natural image prior for variations within the category. We then describe our mechanism for collecting cognitively impenetrable perceptual judgments in Section 3.2.
@@ 3.1 Generating images with varied distortions
We leverage Stable Diffusion v1.4  rombach2022high , which generates diverse and high-quality images that adhere to a given text prompt. We sample images with a prompt of the same category, using the structure ‘‘An image of a <category>’’. The <category> is drawn from image labels in popular datasets: ImageNet deng2009imagenet , CIFAR-10 krizhevsky2009learning , CIFAR-100 krizhevsky2009learning , Oxford 102 Flower nilsback2008automated , Food-101 bossard2014food , and SUN397 xiao2010sun . While the ImageNet labels make up $\sim\!55\%$ of the categories, the superset of categories across these datasets encompass a wide range of objects and scenes.
As illustrated in Figure 1, the resulting images display mid-level variations, such as pose, perspective, and shape, falling in-between pixel/patch-based distortions and categorical differences. Additional triplets and the full list of categories are in the Supplementary Materials (SM). We generate an initial set of 100,000 image triplets, which are later filtered during the labeling process. Each triplet consists of a reference image and two “distortions”, generated from the same prompt.
@@ 3.2 Human perceptual judgments
We strive to gather perceptual judgments that are automatic (requiring little to no cognition), stable (invariant to changes in mental representation), and shared across humans, which requires a cognitively impenetrable judging process. Our main dataset consists of two alternative forced choice (2AFC) judgments on a triplet. Additionally, we collect just noticeable difference (JND) judgments on image pairs, corroborating our findings. Table 1 shows an overview of our dataset, compared to prior perceptual datasets.
@@
To collect cognitively impenetrable triplets, we design our 2AFC experiments to ensure that each triplet included in the dataset obtains a unanimous vote for either $\tilde{x}_{0}$ or $\tilde{x}_{1}$. We divide our experiments into 10 rounds, starting with 100,000 triplets, advancing instances that maintain unanimous votes. Triplets retained through all 10 rounds can earn a maximum of 10 votes, but may have less as a result of the aforementioned sentinels. To maintain dataset quality, we only include triplets with $\geq$6 unanimous judgments in the final dataset.
Ultimately, our 2AFC dataset $\mathcal{D}^{\text{2afc}}=\{(x,\tilde{x}_{0},\tilde{x}_{1}),y\}$ consists of 20,019 triplets with an average of 7 unanimous votes each. We partition our resulting dataset into train, validation, and test components with a random 80/10/10 split. Our dataset is publicly available on our project page.
@@
Intuitively, an image pair that falls below the JND threshold should be more likely to be selected as similar in the 2AFC test; thus a high correlation between 2AFC and JND scores indicates that our judgments are reliable under multiple experimental settings. To assess how well our perceptual tests agree, we collect judgments on the triplets from the 2AFC test set, filtering for triplets that “straddle” the JND threshold. Specifically, given triplet $(x,\tilde{x}_{0},\tilde{x}_{1})\in\mathcal{D}^{\text{2afc}}$, we independently collect JND judgments on pairs $(x,\tilde{x}_{0})$ and $(x,\tilde{x}_{1})$, keeping triplets where humans find one and only one to be identical. In total, our JND dataset contains 411 triplets $\mathcal{D}^{\text{jnd}}=\{(x,\tilde{x}_{0},\tilde{x}_{1}),s\}$, where users chose one of $s\in\{0,1\}$ to be identical.
@@ 4.1 Embeddings as a distance metric
We denote a distance between two images as $D(\cdot,\cdot\hskip 1.42262pt;f_{\theta})$, where $f_{\theta}$ is a feature extractor. We evaluate a variety of state-of-the-art candidate embeddings and metrics. LPIPS zhang2018unreasonable and DISTS ding2020image use CNN backbones, whereas DINO caron2021emerging , CLIP radford2021learning , OpenCLIP Cherti2022ReproducibleSL , and MAE he2021masked use transformer-based backbones dosovitskiy2020image . Following standard practice, distance $D(x,\tilde{x};f_{\theta})=1-\text{cos}\big{(}f_{\theta}(x),f_{\theta}(\tilde{x})\big{)}$ is taken as the cosine distance between the CLS tokens taken from the last layer for DINO and MAE (before and after the layer normalization, respectively), and the embedding vector for CLIP and OpenCLIP. For LPIPS and DISTS, $D(x,\tilde{x};f_{\theta})$ is simply the distance metric itself. Given a triplet $(x,\tilde{x}_{0},\tilde{x}_{1})$, and a feature extractor $f_{\theta}$, the model vote is calculated as:
@@ 4.2 Learning an improved metric
Tuning. Next, we investigate the best method to tune or modify feature extractor $f_{\theta}$. A challenge is the large number of parameters (billions, for ViT backbones), compared to the relatively small amount of perceptual data. Inspired by zhang2018unreasonable , we first try tuning via an MLP head, adding a 1-hidden-layer MLP with a residual connection on top of the pre-trained embeddings. We compare this to fine-tuning through the pre-trained backbones using Low-Rank Adaptation (LoRA) hu2021lora , which we find to achieve better results than full fine-tuning. Using LoRA ($r=16$, dropout $p=0.3$, $\alpha=0.5$) we tune approximately 0.67% of each model’s parameters. In all experiments, LoRA significantly outperformed using the MLP head.
@@ 5.1 How well do existing metrics align with human judgments?
In Figure 5, we show how well metrics align with humans on our dataset, across low-level metrics (PSNR, SSIM), prior learned metrics (LPIPS, DISTS) and recent large vision model embeddings (MAE, CLIP, DINO). The best-performing configurations of existing base models are DINO B/16, MAE B/16, CLIP B/32, and OpenCLIP B/32 (additional settings are in SM Section B.1). However, existing metrics have significant misalignments with humans; for example, DINO has approximately 10% disagreement. To improve performance, we finetune with LoRA (as described in Section 4.2), denoted as Tuned - LoRA, achieving significant alignment improvements over pre-trained baselines (Figure 5) and training an MLP head (Tuned - MLP). Ensembling models by concatenating OpenCLIP, CLIP, and DINO features achieves the highest score across baselines (90.8%), MLP-tuned models (93.4%), and LoRA-tuned models (96.2%). We refer to the best LoRA-tuned ensemble model as DreamSim, and use it for subsequent analysis and applications.
Do different ways of measuring human perception agree? We corroborate our results on our 2AFC test set by evaluating the models against JND perceptual judgments. The high correlation between 2AFC and JND scores indicates that our 2AFC judgments capture a general sense of similarity that can replicated across different settings (Figure 5).
Does improved mid-level perception lead to improved low-level and high-level perception? We use the same models as in Figure 5 to examine how performance on our dataset corresponds to performance on two others: BAPPS, containing image triplets with low-level augmentations, and THINGS, containing triplets with categorical variations. Training on our dataset improves human alignment on BAPPS but decreases alignment on THINGS, suggesting that humans, when making an automatic judgment, tend to focus on appearance-based visual differences (captured in BAPPS and our dataset) rather than high-level object category. We additionally evaluate on the TID2013 tid2013 and KADID-10k kadid10k image quality assessment (IQA) benchmarks as alternative low-level datasets. Similar to BAPPS, we find that our metric outperforms base models and is competitive with low-level metrics despite not training for low-level similarity. For full results, see SM Section B.1.
@@ 5.2 What image attributes affect similarity decisions?
Sensitivity to image ablations. We study which image properties affect our model’s decisions by ablating properties of the image triplets, such as orientation, color, shading, structure, and foreground or background components. To do this, we modify the triplets to change or discard each attribute and compute the agreement between the model decisions on the modified and unmodified triplets. This operation measures the model’s robustness when a given attribute is missing. Note that this analysis aims to provide insight into how our model determines similarity between two images, rather than to compare to human judgments; it is possible that humans may be more sensitive to different image attributes.
@@
As shown in Figure 6, our model is least affected by orientation (tested by flipping the reference image), but ablating color or shading (the L or AB channels in LAB color space), structure  gatys2015texture , or both color and shading together using contours chan2022learning has larger impacts. We conclude that color, shading, and structure contain critical information for our model, but our model can tolerate differences in orientation. To ablate foreground and background components, we segment the image carvekit and replace the corresponding pixels with random uniform noise. This operation removes the color, shading, and texture of the region, even though the same outline remains in both. Removing the background has a smaller impact on model alignment compared to removing the foreground, suggesting that the foreground color and texture is more important for similarity recognition in our model compared to the background.
Figure 7: Alignment with semantic and low-level metrics. We select random triplets from the COCO dataset and assess how well low-level and semantic metrics predict model judgments. Compared to LPIPS & DISTS, our model is more sensitive to objects that appear in individual instances (“things”) and the presence of people, and less sensitive to “stuff” categories, such as sky and road. Compared to OpenCLIP & DINO, color explains more of our model’s decisions. We mark our model’s results with a dashed pink line for comparison.
@@
Alignment with semantic and low-level metrics. Given that some image attributes, like semantics, are difficult to ablate, we complement our ablation analysis by examining how well semantic and handcrafted low-level features align with model judgments. We use 10,000 random image triplets from MS-COCO lin2014microsoft annotated with individual object instances (“things”) and background categories such as grass and sky (“stuff”). We compare our metric to LPIPS, DISTS, DINO, and OpenCLIP (Figure 7). Our model aligns less strongly with RGB-color histogram similarity (calculated using histogram intersection swain1991color ) compared to LPIPS and DISTS, however is more sensitive to color compared to other large vision models, while none align with depth map similarity Eigen2014DepthMP ; Ranftl2020 . Next, we compute alignment with the things-category and stuff-category histograms, which summarize the area occupied by each semantic category. Our model aligns better with things-histogram similarity, whereas other models are comparatively more sensitive to stuff. This result suggests that our model is relatively more sensitive to foreground object instances. Finally, we compute how often metrics align with the per-category intersection of categorical histograms. Our model aligns best with the presence of people (additional results in SM Section B.1).
@@
Figure 8: Nearest-neighbor retrieval across different metrics. Nearest neighbor retrieval on ImageNet-R (top) and COCO (bottom), comparing different metrics. Although the datasets include images outside of our training domain, our model consistently retrieves neighbors with similar appearance and class to that of the query image.